### v2

# Fluctuations and Irreversibility: Historical and Modern Perspectives

Abstract
This article traces the development of fluctuation theory and its deep connection to irreversibility, from equilibrium to near-equilibrium, and finally to far-from-equilibrium systems. Classical fluctuation theorems, which capture the statistical behaviour of thermodynamic systems far from equilibrium, are now well established. Their quantum counterparts, however, remain an active area of research. In this review, we highlight recent advances by linking quantum fluctuation theorems with linear response theory, offering new insights into the nature of quantum fluctuations and irreversibility in the near-equilibrium regime. Particular emphasis is placed on dissipated work in quantum systems as a pathway to observing non-classical effects in quantum thermodynamics. Understanding quantum fluctuations is not only essential for clarifying the foundations of irreversibility but also crucial for the development of novel quantum technologies, including quantum computers, sensors, and metrological devices.
entropy production; fluctuation theorem; linear response theory and quantum thermodynamics
I Introduction
“Thermodynamics is the only physical theory of universal content which I am convinced will never be overthrown.”
— Albert Einstein einstein1926
The control of fire and the mastery of the wheel have been regarded as two of humanity’s greatest inventions. Even today, it is difficult to imagine building any complex structure without the wheel or the underlying principle of circular motion on an axis. However, it was not until 1763 that, while repairing Newcomen engines, James Watt combined the concepts of fire and the wheel to create the steam engine rao2011history . This innovation harnessed the energy of coal flames to produce rotary motion, powering locomotives, factory machinery and setting the stage for the Industrial Revolution. To pursuit the desire to optimise these widely used heat engines, a remarkable group of people such as Sadi Carnot, James Joule, Rudolf Clausius, William Thompson, James Clerk Maxwell and Willard Gibbs laid the foundations of thermodynamics muller2007history . Though initially formulated for equilibrium and reversible processes, thermodynamics eventually came to incorporate the crucial roles of fluctuations and irreversibility — a topic of the present article. From the 1900s onwards, we further unveiled the atomic and molecular world, giving rise to quantum mechanics. More recently, our curiosity has turned to the question: ‘Are quantum thermodynamic machines better than their classical counterparts?’ ghosh2019are Just as the combination of fire and the wheel marked the beginning of the technological revolution, the harnessing of quantum resources may herald a new era. This review aims to briefly trace that remarkable journey.
The organisation of the paper is as follows: Section 2 presents a review of fluctuations in classical equilibrium and near-equilibrium thermodynamics. Section 3 introduces stochastic thermodynamics, where we discuss the various classical fluctuation theorems developed in the 1990s. Extensions to quantum fluctuation relations within the linear response regime are presented in Section 4. Finally, Section 5 concludes the paper.
II Fluctuations in Equilibrium and Near-equilibrium Thermodynamics
“If your theory is found to be against the second law of thermodynamics I can give you no hope.”
— Arthur Eddington eddington1928book
From the middle of the twentieth century, our understanding of thermodynamics has changed dramatically. While classical thermodynamics relies on equilibrium properties, increasing importance is now given towards the study of fluctuations, instabilities, and evolutionary processes at all levels, from chemistry, biology to cosmology kondepudi2015book . Almost all processes in nature are irreversible, in which time-reversal symmetry is broken. The distinction between reversible and irreversible processes was introduced into thermodynamics through the concept of “entropy” — the arrow of time, as Arthur Eddington called it eddington1928book . Unlike Newtonian mechanics, which originated in the study of celestial bodies, equilibrium thermodynamics — as emphasised in the introduction — arose during the Industrial Revolution as a phenomenological theory of macroscopic objects with a much more practical purpose: replacing manual labour with efficient machines by extracting useful work from the motive power of fire carnot1824book . To understand the thermodynamic behaviour of a system, it is therefore imperative to examine how it exchanges energy with its surroundings in the form of heat and work. Thus, the foundation of classical thermodynamics rests on two main pillars: energy and entropy callen1985book . Accordingly, thermodynamic systems are classified into three categories: isolated, closed, and open. In doing so, thermodynamics introduces the notion of temperature — a concept which was absent from classical mechanics. If the temperature of a system is not uniform everywhere, heat will flow until the entire system reaches a single temperature. Such a macroscopic state, characterized by the absence of change on the average, is called an equilibrium state of the system. In particular, traditional thermodynamics focuses on average properties of macroscopic systems and disregards fluctuations. Since thermodynamic systems are typically large, thermal fluctuations are negligible; however, as the system size decreases, both thermal as well as quantum fluctuations become inevitable and must be taken into account binder2019thermodynamics . So, the description of non-equilibrium process, requires not only the averages but also the deviations of various thermodynamic quantities from its equilibrium values.
In its original form, an irreversible transformation was characterized by Clausius clausius1891book in terms of
$N=S-S_{0}-\int\frac{dQ}{T},$ (1)
where $S$ is the entropy of the final state and $S_{0}$ is the entropy of the initial state. According to him, “The magnitude $N$ thus determines the uncompensated transformation (uncompensirte Verwandlung)” clausius1879book . It represents the entropy produced by irreversible processes, which, by virtue of the second law of thermodynamics, can only be positive prigogine1961book . This production of entropy within the system establishes an arrow of time, distinguishing the future from the past. This stands in sharp contrast to the laws of mechanics — both classical and quantum — which are time symmetric and possess no intrinsic time-irreversibility in their fundamental equations. Thus, processes ruled out by the second law of thermodynamics, such as the spontaneous flow of gas molecules from a region of lower concentration to one of higher concentration, don’t contradict with the laws of mechanics kondepudi2015book . This leads to a fundamental question: How can irreversible macroscopic processes, such as the flow of heat from higher to lower temperatures, emerge from the reversible microscopic dynamics of atoms and molecules? The answer lies in the foundations of statistical mechanics, which connect the probabilistic description of microscopic states with the macroscopic behaviour of thermodynamic systems evans2016fundamentals .
To address these pertinent questions, Ludwig Boltzmann was first to introduce an extraordinary relation — engraved on his memorial in the Central Cemetery in Vienna — that entropy is a logarithmic measure of the number of microstates: lebowitz1993boltzmann
$S=k_{B}\ln W.$ (2)
Often, $W$ is referred to as the thermodynamic probability corresponding to the macrostate having entropy, $S$. Now, the random motion of atoms and molecules causes all thermodynamic quantities to fluctuate around their equilibrium values. If the system slightly deviates from its equilibrium value $S_{0}$, we may express entropy as a function of a thermodynamic variable $A_{i}$ (such as internal energy, volume, or number of particles), while keeping all other parameters $A_{j\neq i}$ fixed:
$S=f(A_{i}).$ (3)
Expanding $S$ in a Taylor series around the equilibrium value $A_{i}^{0}$ and retaining terms up to second order, we obtain kondepudi2015book
$S=S_{0}+\bigg(\frac{\partial S}{\partial A_{i}}\bigg)_{A_{i}=A_{i}^{0}}(A_{i}-A_{i}^{0})+\frac{1}{2}\bigg(\frac{\partial^{2}S}{\partial A_{i}^{2}}\bigg)_{A_{i}=A_{i}^{0}}(A_{i}-A_{i}^{0})^{2}+\dots$ (4)
Here, $A_{i}^{0}$ is the equilibrium value of the thermodynamic variable $A_{i}$. Since entropy reaches its maximum value at equilibrium, we have
$\bigg(\frac{\partial S}{\partial A_{i}}\bigg)_{A_{i}=A_{i}^{0}}=0,\qquad\text{and}\qquad\bigg(\frac{\partial^{2}S}{\partial A_{i}^{2}}\bigg)_{A_{i}=A_{i}^{0}}<0.$ (5)
Hence, the deviation of entropy from its equilibrium value due to a fluctuation in $A_{i}$ can be written as
$\Delta S=S-S_{0}=\frac{1}{2}\bigg(\frac{\partial^{2}S}{\partial A_{i}^{2}}\bigg)_{A_{i}=A_{i}^{0}}(A_{i}-A_{i}^{0})^{2},$ (6)
which is strictly negative. Thus, any fluctuation can only reduce the entropy of a system in equilibrium.
Later, Albert Einstein prescribed a formula by rewriting Boltzmann’s idea in reverse: sommerfeld1964book
$W=e^{S/k_{B}},$ (7)
where $S$ is considered to be known empirically, and $W$ is the unknown to be determined using the above relation. From this, we obtain a probability distribution function as a ratio of $W/W_{0}$:
$P(A_{i})=\frac{W}{W_{0}}=e^{\frac{\Delta S}{k_{B}}}=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp[-\frac{(A_{i}-A_{i}^{0})^{2}}{2\sigma^{2}}\bigg],$ (8)
where $W$ and $W_{0}$ are the thermodynamic probabilities corresponding to the entropies $S$ and $S_{0}$, respectively. This is a normalized Gaussian equilibrium probability distribution function, consistent with the central limit theorem, where $\sigma^{2}$ is the variance given by
$\sigma^{2}=\langle(A_{i}-A_{i}^{0})^{2}\rangle=-\frac{k_{B}}{\bigg(\frac{\partial^{2}S}{\partial A_{i}^{2}}\bigg)_{A_{i}=A_{i}^{0}}},$ (9)
and is positive definite kondepudi2015book . As an example, if we calculate the fluctuation in volume for an ideal gas dsray2015note , the relative fluctuation tends to zero for large $N$:
$\frac{\sigma^{2}}{\langle V\rangle^{2}}=\frac{\langle(V-\langle V\rangle)^{2}\rangle}{\langle V\rangle^{2}}=\frac{1}{N}\longrightarrow 0\quad(\text{for }N\to\infty),$ (10)
i.e., equilibrium fluctuations become negligible for a macroscopically large thermodynamic system.
Now, for multiple variables, we introduce the Onsager coordinates $\alpha_{i}=A_{i}-A_{i}^{0}$ (i=1,2,)(i=1,2,...), which represent the fluctuations of different thermodynamic variables around their equilibrium values. Proceeding as before, we expand the entropy, $S=f(\{A_{i}\})$, up to linear-order terms in $\alpha_{i}$s’: kondepudi2015book ; dsray2015note
$S=S_{0}+\sum_{i}\left(\frac{\partial S}{\partial\alpha_{i}}\right)_{\alpha_{i}=0}\alpha_{i}+\frac{1}{2}\sum_{ij}\left(\frac{\partial^{2}S}{\partial\alpha_{i}\partial\alpha_{j}}\right)_{\alpha_{i},\alpha_{j}=0}\alpha_{i}\alpha_{j}+\dots$ (11)
The change in entropy due to fluctuations is then
$\Delta S=S-S_{0}=\frac{1}{2}\sum_{ij}\left(\frac{\partial^{2}S}{\partial\alpha_{i}\partial\alpha_{j}}\right)_{\alpha_{i},\alpha_{j}=0}\alpha_{i}\alpha_{j}$ (12)
Since the system is in equilibrium, this change is negative, as before. So, the equilibrium state is stable against any perturbation that decreases entropy. Conversely, if fluctuation grows, the system is not in equilibrium. This principle is known as Gibbs stability theory kondepudi2015book .
In response to a fluctuation that decreases entropy from its maximum equilibrium value, there will be irreversible processes producing entropy ($\Delta_{i}S$) that spontaneously drive the system back to equilibrium kondepudi2015book . At equilibrium, these processes vanish and 19th-century thermodynamics was largely focused on such idealised reversible transformations. Recognising the connection between entropy and irreversibility, Pierre Duhem began developing a formalism that was later completed by Prigogine prigogine1961book and Onsager onsager1931reciprocal-I ; onsager1931reciprocal-II . In the modern approach, entropy change is calculated in terms of variables that characterise non-equilibrium processes. Even if a system is not in global equilibrium, thermodynamic quantities such as temperature, concentration, pressure, and internal energy remain well-defined locally, i.e., intensive parameters (temperature, pressure, etc.) are well defined within small volume elements, while extensive variables (entropy, internal energy, etc.) are expressed in terms of their corresponding densities. This is the assumption of local equilibrium, which holds pretty well for most physical and chemical systems. prigogine1961book ; kondepudi2015book
Modern formalism due to Prigogine, begins with the infinitesimal change in entropy, written as prigogine1961book ; kondepudi2015book
$dS=d_{e}S+d_{i}S,$ (13)
where $d_{e}S$ is the entropy change due to the exchange of matter and energy with the environment, and $d_{i}S$ is the entropy change due to the uncompensated transformation, i.e., the entropy produced by irreversible transformations. This in turn can be described by the sum of individual thermodynamic forces (such as gradients of temperature, concentration, etc.) and their corresponding thermodynamic fluxes (such as heat flow, diffusion, etc.). The rate of entropy production can thus be expressed as callen1985book ; kondepudi2015book
$\dot{\Sigma}=\frac{d_{i}S}{dt}=\sum_{k}F_{k}J_{k},$ (14)
where $F_{k}$ and $J_{k}$ are the respective thermodynamic forces and fluxes. The entropy production rate and the entropy production itself are always positive ($\dot{\Sigma}>0,\;d_{i}S>0$). This statement is stronger and more general than the classical formulation that the entropy of an isolated system can only increase. This insight by Prigogine earned him the Nobel Prize in Chemistry in 1977 prigogine1977time .
To obtain Eq. (14) for the $\dot{\Sigma}$ associated with fluctuations in system variables, the thermodynamic force can be written as, $F_{k}=\frac{\partial\Delta_{i}S}{\partial\alpha_{k}}$, conjugate to the thermodynamic flux $J_{k}=\frac{\partial\alpha_{k}}{\partial t}$. If the deviations of the forces from their equilibrium values are small, the fluxes can be expressed as linear functions of the forces, $J_{k}=\sum_{j}L_{kj}F_{j}$, where $L_{kj}$ are constants known as kinetic coefficients. Thus, the rate of entropy production can be calculated as kondepudi2015book
$\dot{\Sigma}=\frac{\partial\Delta_{i}S}{\partial t}$ $=$ $\sum_{k}\frac{\partial\Delta_{i}S}{\partial\alpha_{k}}\frac{\partial\alpha_{k}}{\partial t}$ (15)
$=$ $\sum_{k}F_{k}J_{k}=\sum_{jk}F_{j}L_{jk}F_{k}>0.$
Here, $\mathbf{L}$ is a positive definite matrix whose elements satisfy the symmetry condition $L_{kj}=L_{jk}$. These reciprocal relations are linked to thermodynamic cross-effects and were first observed by Lord Kelvin and others in the study of thermoelectric phenomena (such as the Seebeck (1821) and Peltier (1834) effects in metal wires) kondepudi2015book ; gupt2024graph . At that time, they were regarded merely as conjectures. A rigorous mathematical derivation was later provided by Lars Onsager in 1931, after which they became known as the Onsager reciprocal relations onsager1931reciprocal-I ; onsager1931reciprocal-II . The foundation of Onsager’s theory rests on the properties of steady-state fluctuations:
$\langle\alpha_{j}(t)\alpha_{k}(t+\tau)\rangle=\langle\alpha_{j}(t)\alpha_{k}(t-\tau)\rangle=\langle\alpha_{j}(t+\tau)\alpha_{k}(t)\rangle.$ (16)
The first equality reflects time-reversal symmetry, while the second expresses time-translation invariance. Taken together, these properties — often referred to as microscopic reversibility or the principle of detailed balance — lead directly to the reciprocal relations $L_{kj}=L_{jk}$. The importance of this result is so profound that it is regarded as “fourth law of thermodynamics,” deffner2019quantumthermo a contribution for which Onsager was awarded the Nobel Prize in Chemistry in 1968 onsager1968the .
We conclude this section by highlighting the deep connection between entropy and fluctuations in equilibrium and near-equilibrium regimes. This connection explains why equilibrium thermodynamic properties are primarily governed by average values and how microscopic fluctuation symmetries give rise to thermodynamic cross-effects at the macroscopic scale. To fully comprehend how irreversibility emerges at macroscopic length scales from time-reversible microscopic dynamics, one must turn to the field of stochastic thermodynamics evans2016fundamentals , which came to the forefront in the 1990s.
III Emergence of Stochastic Thermodynamics and Fluctuation Theorems
“If we shift our focus away from equilibrium states, we find a rich universe of nonequilibrium behavior.”
— Christopher Jarzynski jarzynski2015diverse
Over the past three decades, stochastic thermodynamics has become a leading framework for understanding diverse non-equilibrium phenomena, ranging from molecular motors and cellular reactions to colloids and nanoscale systems udo2012stochastic ; peliti2021stochastic ; limmer2024statistical ; seifert2025stochastic . Often, thermodynamics on this scale refers as nanoscale thermodynamics broedersz2022twenty , and researchers are now attempting to extend this concept to the quantum realm, which will be discussed in the following section benenti2017fundamental . Broadly speaking, stochastic thermodynamics describes mesoscopic nonequilibrium systems in contact with equilibrium heat bath. Here, the system interacts randomly with the reservoir, so that at time $t$ it may be found in a state $x$ with probability $P(x;t)$. In such small systems, fluctuations play a pivotal role in controlling the dynamics of the system, and thermodynamic quantities such as heat, work, and entropy themselves fluctuate. Unlike in the classical case, these quantities don’t take on single fixed values but are instead described by probability distributions, with mean values serving as the relevant averages. Remarkably, these distributions for thermodynamic observables fulfil some symmetry relations known as fluctuation theorems (FTs) seifert2025stochastic . Here, we will discuss a few such relations.
III.0.1 Evans-Searles fluctuation theorem (1994):
Evans and Searles proposed the first fluctuation relation related to general dissipation in simulations of sheared fluids evans1994equilibrium . Here, we present a general version of the theorem within Hamiltonian dynamics. Consider an isolated Hamiltonian system of interacting particles. The state of the system is represented by a phase-space vector containing the coordinates and canonical momenta of all particles, in an Avogadro-dimensional space, denoted as $\Gamma\equiv\{\textbf{q}_{i},\textbf{p}_{i}\}$. The Hamilton’s equations for such an autonomous system are given by
$\dot{\textbf{q}}_{i}$ $=$ $\frac{\partial H(\{\textbf{q}_{i},\textbf{p}_{i}\})}{\partial\textbf{p}_{i}},\qquad\dot{\textbf{p}}_{i}=-\frac{\partial H(\{\textbf{q}_{i},\textbf{p}_{i}\})}{\partial\textbf{q}_{i}}.$ (17)
Due to the large number of degrees of freedom, it is not feasible to specify the initial condition of the system by explicitly providing all $\{\mathbf{q}_{i}(0),\mathbf{p}_{i}(0)\}$. Instead, we introduce the probability distribution function (PDF), $f(\Gamma,t)$, with the initial condition
$f(\Gamma,0)=\frac{e^{-F(\Gamma)}}{\int_{D}d\Gamma\,e^{-F(\Gamma)}}.$ (18)
Here, $F(\Gamma)$ is an arbitrary single-valued real function defined over a specified phase-space domain $D$. Defining the PDF as $f(\Gamma,t)$, the probability of finding the system in an infinitesimal volume of phase-space $d\Gamma_{t}$ at time $t$ is given by $P(\Gamma_{t},t)=f(\Gamma_{t},t)\,d\Gamma_{t}$. Using the total time derivative of $f(\Gamma_{t},t)$ and the continuity equation for probability flow in phase space, we obtain
$\frac{df}{dt}=\frac{\partial f}{\partial t}+\frac{\partial f}{\partial\Gamma}\dot{\Gamma},\qquad\text{and}\qquad\frac{\partial(f\dot{\Gamma})}{\partial\Gamma}+\frac{\partial f}{\partial t}=0.$ (19)
respectively. From Eq. (19), the time evolution of $f(\Gamma_{t},t)$ can be calculated as follows:
$\frac{df}{dt}=-f\Lambda,\qquad\text{where}\quad\Lambda=\frac{\partial\dot{\Gamma}}{\partial\Gamma}.$ (20)
Integrating Eq. (20) from $t=0$ to $t=\tau$, yields
$f(\Gamma_{\tau},\tau)=f(\Gamma_{0},0)\,e^{-\int_{0}^{\tau}\Lambda\,dt}.$ (21)
Since the probability $P(\Gamma_{t},t)$ of a specific phase-space volume is conserved during evolution, we have
$f(\Gamma_{\tau},\tau)\,d\Gamma_{\tau}=f(\Gamma_{0},0)\,d\Gamma_{0}.$ (22)
Combining Eqs. (21)-(22), we find that the compression factor for the infinitesimal phase-space volume is given by
$\frac{d\Gamma_{\tau}}{d\Gamma_{0}}=e^{\int_{0}^{\tau}\Lambda\,dt},$ (23)
which plays a central role in formulating FTs.
To understand the origin of the time arrow, we first note that the microscopic equations of motion, Eq. (17), is invariant under time reversal. Since the dynamics here are deterministic, any trajectory is uniquely determined by its initial phase-space point. Let $S^{\tau}$ denote the time evolution operator that acts over the time interval 0 to $\tau$. Thus, for a trajectory starting at $\Gamma_{0}$, we have $\Gamma_{\tau}=S^{\tau}\Gamma_{0}$. Now, let $M^{\tau}$ be the time-reversal mapping operator, which reverses the momenta as $M^{\tau}\Gamma=M^{\tau}\{\textbf{q}_{i},\textbf{p}_{i}\}=\{\textbf{q}_{i},-\textbf{p}_{i}\}$. So, for the “anti-trajectory” conjugate to $\Gamma_{0}$, we can depict the initial phase-space point by $\Gamma^{*}_{0}=M^{\tau}\Gamma_{\tau}=M^{\tau}S^{\tau}\Gamma_{0}$. Since time reversal leaves the phase-space volume unchanged, we have $d\Gamma_{0}^{*}=d\Gamma_{\tau}$, which, upon using Eq. (23), becomes
$d\Gamma_{0}^{*}=d\Gamma_{0}\,e^{\int_{0}^{\tau}\Lambda dt}.$ (24)
Next, instead of a single trajectory, we consider a bundle of trajectories occupying a phase-space volume $d\Gamma_{t}$ with probability $P(d\Gamma_{t},t)$. Microscopic reversibility requires equal probabilities for both forward and reverse bundles of trajectories, i.e., $P(d\Gamma_{0},0)=P(d\Gamma_{0}^{*},0)$, which leads to
$f(d\Gamma_{0},0)\,d\Gamma_{0}=f(d\Gamma_{0}^{*},0)\,d\Gamma_{0}^{*}.$ (25)
Since it is based on the principle of microscopic reversibility, Eq. (25) could be regarded as the criterion for equilibrium. Combining it with Eq. (24) yields
$\frac{f(d\Gamma_{0},0)}{f(d\Gamma_{0}^{*},0)}=e^{\int_{0}^{\tau}\Lambda dt}.$ (26)
If the system is not in equilibrium, then $P(d\Gamma_{0},0)\neq P(d\Gamma_{0}^{*},0)$. Hence, we can, in general, define a dissipation function along a trajectory originating from the phase-space point $\Gamma_{0}$ as
$\Omega_{\tau}(\Gamma_{0})=\ln\bigg(\frac{P(d\Gamma_{0},0)}{P(d\Gamma_{0}^{*},0)}\bigg.)=\ln\bigg(\frac{f(d\Gamma_{0},0)}{f(d\Gamma_{0}^{*},0)}\bigg.)-\int^{\tau}_{0}\Lambda dt,$ (27)
which quantifies the deviation from equilibrium. For $\Omega_{\tau}(\Gamma_{0})$ to be well defined over the phase space domain $D$, the system must be ergodically consistent over $D$ evans2016fundamentals . It is evident that $\Omega_{\tau}(\Gamma_{0})=0$ at equilibrium. Since $\Omega_{\tau}(\Gamma_{0})$ characterizes the reversibility of a set of trajectories, we can evaluate the relative probability of the dissipation function taking opposite values:
$\frac{\mathcal{P}(\Omega_{\tau}=A)}{\mathcal{P}(\Omega_{\tau}=-A)}=\frac{\int d\Gamma_{0}\delta(\Omega_{\tau}(\Gamma_{0})-A)f(\Gamma_{0},0)}{\int d\Gamma_{0}^{*}\delta(\Omega_{\tau}(\Gamma_{0}^{*})+A)f(\Gamma_{0}^{*},0)}=e^{A},$ (28)
where we used the key property of the odd parity of the dissipation function, i.e., $\Omega_{\tau}(\Gamma_{0})=-\Omega_{\tau}(\Gamma_{0}^{*})$, which follows directly from its definition in Eq. (27). Equation (28) is known as the Evans–Searles fluctuation theorem (ESFT) evans1994equilibrium . This implies that dissipation functions taking positive values for forward irreversible trajectories are more probable than for the corresponding reverse antitrajectories. Thus, the dissipation function can be viewed as a measure of the temporal asymmetry inherent in bundles of trajectories originating from an initial distribution of states. In defining the dissipation function in Eq. (27), which characterizes deviations from equilibrium, we did not restrict ourselves to the near-equilibrium situation. As a consequence, the ESFT derived here applies to nonequilibrium fluctuations arbitrarily far from equilibrium and to systems of any size, without requiring the classical thermodynamic limit. Most importantly, the ESFT demonstrates that even deterministic, time-reversible equations of motion can give rise to irreversibility evans2016fundamentals . The ESFT has also been verified experimentally, for instance by Wang et al. (2002) wang2002experimental and Reid et al. (2004) reid2004reversibility .
One can further evaluate the phase-space average of the dissipation function as $\langle\exp(-\Omega_{\tau})\rangle=1$. Since $\exp(-\Omega_{\tau})$ is a convex function, Jensen’s inequality gives seifert2025stochastic
$\exp(-\langle\Omega_{\tau}\rangle)\leq\langle\exp(-\Omega_{\tau})\rangle=1,$ (29)
which implies $\langle\Omega_{\tau}\rangle\geq 0$. This shows that although the dissipation function defined in Eq. (27) can take positive, negative, or zero values, its average is always nonnegative. This is the statement of the second law of thermodynamics, highlighting the link between microscopic reversible dynamics and macroscopic irreversible behavior, and providing a resolution of the Loschmidt paradox evans2016fundamentals . In the following, we extend the FTs to more practical thermodynamic quantities.
III.0.2 Jarzynski equality (1997):
Jarzynski developed a more practical relation that allows us to relate the work $W$ and the free energy change $\Delta F$ for a non-quasistatic process between two sates of equilibrium. It is an equality and hence contains more information than classical thermodynamics, which only connects them through an inequality $W\geq\Delta F$ jarzynski1997nonequilibrium ; jarzynski2010equalities .
Consider a driven isolated system with a $\lambda$-dependent Hamiltonian $H(\Gamma,\lambda)$ defined on the domain $D_{\lambda}$ of the phase space. At equilibrium, for each value of $\lambda$ and inverse temperature $\beta$, there corresponds a free energy
[…]
The system is driven through a regular protocol $\lambda(t)$, with $0\leq t\leq\tau$, from an initial value $\lambda(0)$ to a final value $\lambda(\tau)$. For an initial phase point $\Gamma_{0}$, the work applied up to time $\tau$ is defined as
$\begin{split}W(\Gamma_{0})&=H(\Gamma_{\tau}(\Gamma_{0}),\lambda(\tau))-H(\Gamma_{0},\lambda(0))\\ &=\int_{0}^{\tau}dt\partial_{\lambda}H(\Gamma_{t},\lambda)\bigg|_{\lambda=\lambda(t)}\partial_{t}\lambda(t).\end{split}$ (31)
If the system is initially thermalised at inverse temperature $\beta$, the normalised initial canonical distribution $p_{i}(\Gamma_{0})$ is given by
$p_{i}(\Gamma_{0})=e^{-\beta[H(\Gamma_{0},\lambda(0))-F(\beta,\lambda(0))]}$ (32)
Now, for any auxiliary final distribution $p_{f}(\Gamma_{1})$, normalised on $D_{1}$, we can write
$p_{f}(\Gamma_{1})=p_{f}(\Gamma_{\tau}(\Gamma_{0}))=e^{-\beta[H(\Gamma_{\tau}(\Gamma_{0})),\lambda(\tau))-F(\beta,\lambda(\tau)]}$ (33)
Since, for a regular protocol, the dynamical image of $D_{0}$ after time $\tau$ becomes $D_{1}$, the normalisation of $p_{f}(\Gamma_{1})$ is guaranteed. So, we can write seifert2025stochastic
$\begin{split}\sum_{\Gamma_{1}\in D_{1}}p_{f}(\Gamma_{1})=1=&\sum_{\Gamma_{0}\in D_{0}}p_{i}(\Gamma_{0})\frac{p_{f}(\Gamma_{\tau}(\Gamma_{0}))}{p_{i}(\Gamma_{0})}\\ =&~\bigg\langle\frac{p_{f}(\Gamma_{\tau}(\Gamma_{0}))}{p_{i}(\Gamma_{0})}\bigg\rangle=\langle e^{-\Omega(\Gamma_{0})}\rangle\end{split}$ (34)
This remarkable property, which applies for the distribution of work and several other thermodynamic quantities, is known as the integral fluctuation theorem (IFT) udo2012stochastic ; seifert2025stochastic . From Eqs. (32)-(34), we identify
$\begin{split}\Omega(\Gamma_{0})=&\beta[H(\Gamma_{\tau}(\Gamma_{0}),\lambda(\tau))-H(\Gamma_{0},\lambda(0))-\Delta F]\\ =&\beta[W(\Gamma_{0})-\Delta F]\end{split}$ (35)
with $\Delta F=F(\beta,\lambda(\tau))-F(\beta,\lambda(0))$. Substituting Eq. (35) into Eq. (34), the corresponding IFT reduces to the celebrated Jarzynski equality (JE) jarzynski1997nonequilibrium :
$\langle\exp(-\beta W(\Gamma_{0}))\rangle=\exp(-\beta\Delta F).$ (36)
This equality relates an average over the work required in this process to the free-energy difference between the two canonical equilibrium states: one corresponding to the value $\lambda(\tau)$ of the control parameter at time $\tau$ and other to its initial value $\lambda(0)$, both taken w.r.t. the same inverse temperature $\beta$. Further, applying Jensen’s inequality to Eq. (36), gives $\mathcal{W}=\langle W(\Gamma_{0})\rangle\geq\Delta F$ udo2012stochastic . If we define the dissipative work as $W_{\text{diss}}=W-\Delta F$, this inequality can be rewritten as $\langle W_{\text{diss}}\rangle\geq\ 0$ seifert2025stochastic . For smaller systems, unlike in the thermodynamic limit, $W_{\text{diss}}$ need not always be positive for individual realisations of the process; it may even be negative. Nevertheless, upon averaging, it becomes consistent with classical thermodynamics. If we identify $W_{\text{diss}}=T\Sigma$, where $\Sigma$ is the entropy production, then combining the IFT with Jensen’s inequality leads to $\langle\Sigma(\Gamma)\rangle\geq 0$ which is recognised as the second law of stochastic thermodynamics peliti2021stochastic . The application of JE to mesoscopic systems in modern times has been remarkable. Notably, JE places no restriction on the speed of the transformation and thus remains valid even far from equilibrium. This makes it a powerful tool for determining free-energy differences in molecular dynamics simulations, driven harmonic oscillators, and even real biomolecular experiments broedersz2022twenty . Many researchers view JE as having brought nineteenth-century thermodynamics into the twenty-first century, given its profound implications deffner2019quantumthermo .
III.0.3 Crooks Theorem (1999):
Although formulated after Jarzynski’s work, Crooks’ theorem provides a more general framework, with the JE arising as a direct consequence. It also illustrates the origin of the arrow of time by introducing the distribution of work: $\mathcal{P}_{F}(W)$ for the forward process, governed by the protocol $\lambda(t)$, and $\mathcal{P}_{R}(-W)$ for the reverse process, governed by the protocol $\lambda(\tau-t)$. Both processes start from thermal equilibrium at inverse temperature $\beta$. Accordingly, we obtain
$\begin{split}&\mathcal{P}_{R}(-W)=\\ =&\sum_{\Gamma^{R}_{0}\in D_{R}}\delta(W_{R}(\Gamma^{R}_{0})+W)e^{-\beta[H(\Gamma^{R}_{0},\lambda(\tau))-F(\beta,\lambda(\tau))]}\\ =&\sum_{\Gamma_{0}\in D_{F}}\delta(-W_{F}(\Gamma_{0})+W)e^{-\beta[H(\Gamma_{0},\lambda(0))+W_{F}(\Gamma_{0})-F(\beta,\lambda(\tau))]}\\ =&\mathcal{P}_{F}(W)e^{-\beta(W-\Delta F)}\end{split}$ (37)
Finally, we can write crooks1998nonequilibrium ; crooks1999entropy
$\frac{\mathcal{P}_{F}(W)}{\mathcal{P}_{R}(-W)}=e^{\beta(W-\Delta F)}.$ (38)
This is the renowned Crooks fluctuation theorem (CFT), which shows that the reverse process is exponentially suppressed, giving rise to the direction of time, much like the second law of thermodynamics. In the derivation, we use the antisymmetry of work-energy conservation and the fact that momenta appear quadratically in the Hamiltonian, i.e.,
$\begin{split}H(\Gamma_{0}^{R},\lambda(\tau))=&H(M^{\tau}\Gamma_{\tau},\lambda(\tau))=H(\Gamma_{\tau},\lambda(\tau))\\ =&H(\Gamma_{0},\lambda(0))+W(\Gamma_{0}).\end{split}$ (39)
The CFT relation in Eq. (38) is stronger than JE [cf. Eq. (36)], since JE can be derived by integrating both sides of Eq. (38) and using $\int dW\mathcal{P}_{R}(-W)=1$. Collin et al. (2005) collin2005verification experimentally verified Crooks’ theorem in a single-molecule RNA hairpin folding–unfolding experiment using optical tweezers. In this experiment, $W_{\text{diss}}$ was sometimes negative, but on average it was positive. Thus, for small biological systems, this theorem provides a clear-cut framework for calculating free-energy changes from work distributions.
III.0.4 The different DFT and furthermore:
If we consider symmetric driving protocol, $\lambda(t)=\lambda(\tau-t)$, then the work distributions for the forward and backward processes become identical and $\Delta F=0$. In this case, the Crooks theorem reduces to
$\frac{\mathcal{P}(W)}{\mathcal{P}(-W)}=e^{\beta W},$ (40)
which is an example of the detailed fluctuation theorem (DFT) seifert2025stochastic . The DFT is a stronger result than the IFT. While the IFT provides a single global constraint on the work distribution (and more generally on entropy production), the DFT establishes a local balance relation between the probabilities of observing a trajectory and its corresponding antitrajectory. Consequently, in a steady-state system, it is sufficient to know $\mathcal{P}(W)$ for positive values of $W$ in order to reconstruct the entire distribution peliti2021stochastic . Both the JE and the CFT can be regarded as special cases of the IFT and DFT when the initial and final states are equilibrium states.
There exists a vast literature on FTs, depending on the observables of interest, the underlying dynamics of the system, and the hypotheses or models considered. Evans and Searles evans1994equilibrium first proposed such a relation. Subsequently, Gallavotti and Cohen gallavotti1995dynamical derived a detailed fluctuation theorem for the entropy production rate in deterministic chaotic systems. Later, the JE jarzynski1997nonequilibrium and CFT crooks1999entropy were developed for systems driven between two equilibrium states, becoming landmark results that enabled equilibrium free-energy differences to be estimated from non-equilibrium measurements. This also led to the rediscovery of earlier works by Bochkov and Kuzovlev bochkov1977general , although their definition of work differs from the present one.
Moreover, introduction of trajectory probabilities for forward and backward processes naturally gave rise to a variety of FTs. For entropy production, if we separate its contributions into two parts — one arising from time dependence and another from external driving (adiabatic and non-adiabatic entropy production) — the IFT leads to the Hatano-Sasa theorem hatano2001steady , which constitutes a cornerstone in the generalisation of FDR (fluctuation-dissipation relation) to non-equilibrium steady state. Subsequently, Rao and Esposito rao2018detailed developed a unified framework that systematically derives many of the FTs reported in the literature. In addition to work and entropy, considerable attention has been devoted to the fluctuation symmetry of heat baiesi2006fluctuation ; noh2012fluctuation . Jarzynski and Wójcik jarzynski2004classical analysed the statistics of heat exchange between two systems coupled to distinct thermal baths, deriving a fluctuation relation that is now widely referred as exchange fluctuation theorem (XFT/EFT) timpanaro2019thermodynamic . More recently, the study of joint statistics of thermodynamic quantities and their associated FTs has attracted considerable attention. Interestingly, while the individual quantities may or may not obey an FT on their own, their joint probability distribution often satisfies a joint fluctuation theorem, thereby revealing richer symmetry structures in far from equilibrium thermodynamics miller2021joint ; chen2023hierarchical .
So far, we have provided an account of the development of FTs in a classical settings. To understand their quantum counterparts, however, we must dig deeper. The key difference arises because any measurement of quantum states or observables disturbs the system and its trajectories, unlike in classical systems. In fact, continuous monitoring can even freeze the dynamics due to the quantum Zeno effect aulettaparisi2009quantum . In the quantum regime, the very definitions of work, heat, and entropy must be handled with care. Thus, exploring quantum versions of FTs is a rich and promising area of current research.
IV Linear Response Theory and Quantum Work Fluctuations
“The idea of dissipation of energy depends on the extent of our knowledge.”
— James Clerk Maxwell maxwell1890the
Up to this point, we have discussed the role of fluctuations in a thermodynamic setup, where general thermodynamic quantities vary due to interactions with heat bath. In 1957, Ryogo Kubo kubo1957statistical laid a concrete foundation within the Hamiltonian framework, showing how any arbitrary dynamical observable of a system coupled to a thermal bath deviates from equilibrium under an external drive, and how irreversible macroscopic behavior appears as a direct consequence of microscopic fluctuations kubo1966thefluctuation when observed through the lens of linear response of the system. Since then, Linear Response Theory (LRT) kubo1983book ; zwanzig2001nonequilibrium ; pottier2010nonequilibrium ; balakrishnan2020elements has held a central place as a powerful tool for studying near-equilibrium systems, with far-reaching applications in fluid dynamics, transport phenomena (both classical and quantum), condensed matter physics, many-body problems, chemical kinetics, and beyond nitzan2024chemical . More recently, it started elucidating its impact on quantum thermodynamics (QTD) guarnieri2024generalized ; gupt2024graph .
IV.0.1 Introduction to LRT:
Here, we establish the Kubo formulas in parallel for both classical balakrishnan2020elements ; balakrishnan2020mathematical ; kubo1983book ; pottier2010nonequilibrium ; zwanzig2001nonequilibrium and quantum mechanics kubo1983book ; pottier2010nonequilibrium . These formulas describe correlation of fluctuations, such as response and relaxation functions between observables of different dynamical variables. We discuss the properties and applications of these response and relaxation functions in the context of QTD.
Consider a physical system (more precisely, a subsystem of interest) in contact with a heat bath at inverse temperature $\beta=1/k_{B}T$. Initially, the system is governed by a time-independent Hamiltonian $H_{0}(\{\mathbf{q}_{i},\mathbf{p}_{i}\})$, which describes its natural dynamics. The normalized canonical equilibrium phase-space density (or, equilibrium density operator in case of quantum mechanics) is given by
$\rho_{\text{eq}}=\frac{e^{-\beta H_{0}}}{\mathrm{Tr}[e^{-\beta H_{0}}]}.$ (41)
The equilibrium average of an arbitrary observable $B(q,p)$ is then
$\langle B\rangle_{\rho_{\text{eq}}}=\frac{\mathrm{Tr}[\rho_{\text{eq}}B]}{\mathrm{Tr}[\rho_{\text{eq}}]}=\frac{\int dq\int dp\,B(q,p)\,e^{-\beta H_{0}(q,p)}}{\int dq\int dp\,e^{-\beta H_{0}(q,p)}}.$ (42)
We now perturb the system, starting from an initial time $t_{0}$ (eventually we take $t_{0}\to$-$\infty$) by applying a time-dependent external force $\lambda(t)$. This force is spatially homogeneous and couples with an arbitrary observable $A$ of the system. The perturbed Hamiltonian is then $H(t)=H_{0}+H^{\prime}(t)=H_{0}-\lambda(t)A$. For $t\geq t_{0}$, the density operator (or the phase-space distribution in the classical case), $\rho(t)$ evolves according to the Liouville–von Neumann equation breuer2002book ; nitzan2024chemical
$\frac{d\rho(t)}{dt}=-i\,\mathcal{L}\,\rho(t),$ (43)
where $\mathcal{L}$ is the Liouville superoperator, defined as
$\mathcal{L}X=\begin{cases}i\{H,X\},&\text{(classical)},\\[8.0pt] \displaystyle\frac{1}{\hbar}[H,X],&\text{(quantum)}.\end{cases}$ (44)
Here, $\{\,\cdot\,$$\cdot\,\}$ and $\cdot\,$$\cdot\,$] denote the Poisson bracket and the commutator, respectively nitzan2024chemical . This formulation allows us to treat classical and quantum dynamics in a unified manner. With this choice one recovers the standard forms:
$\frac{d\rho}{dt}=\{H,\rho\}\quad\text{(classical)},\quad\frac{d\rho}{dt}=-\frac{i}{\hbar}[H,\rho]\quad\text{(quantum)}.$ (45)
For any dynamical observable $\mathcal{O}$ (no explicit time dependence),
$\frac{d\mathcal{O}}{dt}=i\,\mathcal{L}\,\mathcal{O},$ (46)
so the formal time evolution is
$\mathcal{O}(t)=e^{\,i\mathcal{L}(t-t_{0})}\mathcal{O}(t_{0});\;\qquad\rho(t)=e^{-\,i\mathcal{L}(t-t_{0})}\rho(t_{0}).$ (47)
Since the Liouvillian is Hermitian (one can show $\mathcal{L}=\mathcal{L}^{\dagger}$), the operator $e^{\pm i\mathcal{L}t}$ serves as the time evolution operator kubo1983book . We now investigate the evolution of the density operator under an external perturbation in the linear regime. The Liouvillian can be split as $\mathcal{L}=\mathcal{L}_{0}+\mathcal{L}_{\text{pert}}$, associated with $H_{0}$ and $H^{\prime}$ respectively. The perturbed density operator is written as $\rho(t)=\rho_{\text{eq}}+\delta\rho(t)$. Inserting these into Eq. (43), we obtain balakrishnan2020elements ; zwanzig2001nonequilibrium
$\frac{d\delta\rho(t)}{dt}=-i\mathcal{L}_{\text{pert}}\rho_{\text{eq}}-i\mathcal{L}_{0}\delta\rho(t),$ (48)
where we neglect the higher-order term $i\mathcal{L}_{\text{pert}}\delta\rho(t)$. Also, we set $\frac{d\rho_{\text{eq}}}{dt}=0$ and $i\mathcal{L}_{0}\rho_{\text{eq}}=0$, since $\rho_{\text{eq}}$ is stationary. Solving this equation, we obtain
$\delta\rho(t)=\int_{-\infty}^{t}\lambda(t^{\prime})\,\mathcal{C}[A(t^{\prime}-t),\rho_{eq}]\,dt^{\prime},$ (49)
where the operation $\mathcal{C}[\cdot,\cdot]$ is defined as
$\mathcal{C}[X,Y]=\begin{cases}\dfrac{i}{\hbar}[X,Y],&\text{(quantum)},\\[10.00002pt] \{X,Y\},&\text{(classical)}.\end{cases}$ (50)
[…]
$\langle B\rangle_{\delta\rho(t)}=\operatorname{Tr}\{\delta\rho(t)\,B(0)\}=\int_{-\infty}^{t}dt^{\prime}\lambda(t^{\prime})\phi_{AB}(t-t^{\prime})$ (51)
where, the response function is defined as kubo1957statistical ; kubo1983book
$\phi_{AB}(t-t^{\prime}):=\frac{1}{i\hbar}\operatorname{Tr}\!\left\{\rho_{\rm eq}[A(t^{\prime}),B(t)]\right\},$ (52)
and depends only on the equilibrium state. This response of the observable $B$ due to external drive $\lambda(t)$ through the observable $A$, is linear, causal, and retarded
 balakrishnan2020mathematical .
For a relaxation measurement, the system is first driven by an external force from $t=$-$\infty$ to $t=0$. When the force is suddenly switched off, any perturbed observable $B$ relaxes back to equilibrium as
$\begin{split}\langle B\rangle_{\delta\rho(t)}&=\int_{-\infty}^{0}dt^{\prime}\,\lambda(t^{\prime})\,\phi_{AB}(t-t^{\prime})\\ &=\lambda\int_{-\infty}^{t}ds\,\phi_{AB}(s),\quad t>0.\end{split}$ (53)
The relaxation function is then defined as kubo1957statistical ; kubo1983book
$\Phi_{AB}(t)=\lim_{\epsilon\to 0}\int_{-\infty}^{t}\phi_{AB}(s)e^{-\epsilon s}\,ds,$
which characterizes how $B$ relaxes after the perturbation is removed. Using Kubo’s identity for the canonical ensemble nitzan2024chemical ; bellac2004equilibrium ,
$A,\rho_{\rm eq}]=e^{-\beta H_{0}}\int_{0}^{\beta}e^{-\alpha H_{0}}[H_{0},A]e^{-\alpha H_{0}}\,d\alpha,$
(where $\alpha$ is a dummy variable), the relaxation function can be written as
$\begin{split}\Phi_{AB}(t-t^{\prime})&=\int_{0}^{\beta}\operatorname{Tr}\!\big(\rho_{\rm eq}A(-i\hbar\alpha)B(t-t^{\prime})\big)\,d\alpha\\ &\quad-\int_{0}^{\beta}\operatorname{Tr}(\rho_{\rm eq}A)\operatorname{Tr}(\rho_{\rm eq}B)\,d\alpha\end{split}$ (54)
This function is also referred to as the canonical correlation function. However, in quantum mechanics, it is often more useful to express the relaxation function in a form that accounts for the non-commutativity of operators, leading to the symmetrised correlation function, denoted as $\Psi_{AB}(t-t^{\prime})$ kubo1983book . Since it accurately captures the response (i.e., the fluctuation from equilibrium) caused by an external disturbance through time correlations between two dynamical observables, LRT has become increasingly important in quantum thermodynamics, where the control gelbwaser2015thermodynamics and optimisation of quantum devices governed by fluctuations das2024fluctuations , particularly quantum fluctuations, are of growing interest. In our context, the autocorrelation function ($A=B$), which measures how an observable correlates with itself at different times, is of central importance and will be discussed in the next section.
IV.0.2 Quanum Work Fluctuations and LRT:
We live in an era where we aim to reach, utilise, and design systems on scales that have never been achieved before rossnagel2016single ; myers2022quantum ; gupt2022PRE ; sur2023quantum . To fully understand and optimise machines at the quantum level, where thermal and quantum fluctuations play equally important roles, we need to deepen our theoretical understanding of FTs, a.k.a., quantum fluctuation theorems (QFTs) hanggitalkner2015theotherqft ; funo2018quantumft ; strasberg2022quantum ; potts2024quantumthermodynamics . While QFTs can be formulated for various thermodynamic quantities, in this article we focus exclusively on work. Within the LRT framework, formulating QFTs for dissipative work has yielded new insights andrieux2008quantum . More recently, Guarnieri et al. guarnieri2024generalized demonstrated that even near equilibrium, the work distribution is strongly influenced by quantum fluctuations, which are captured through the relaxation functions.
For example, consider a quantum system with Hamiltonian $H_{0}$, initially thermalised in the normalised Gibbs state $\rho_{\text{eq}}=e^{-\beta H_{0}}$. The system is then perturbed unitarily as $H_{t}=H_{0}-\lambda(t)A$ over $t\in[0,\tau]$, where $A$ is an observable and $\lambda(t)$ is a weak dimensionless driving protocol ($\lambda(t)\ll 1$, $\|A\|=1$). At $t=\tau$, when the drive is switched off, the relaxation of $A(\tau)$ is captured by the symmetrized autocorrelation function $\Psi_{0}(=\Psi_{AA})(t-t^{\prime})$, which encodes correlations of $A$ at different times. Since changes in $A$ also modify the Hamiltonian, $\Psi_{0}$ links correlations between $H_{\tau}$ and $H_{0}$, providing a natural connection to the very definition of quantum work guarnieri2024generalized ; talkner2007fluctuation .
The most accomplished and widely used definition of work in QTD is the two-time projective measurement strasberg2022quantum ; deffner2019quantumthermo (TPM) scheme, where the system’s energy is measured at the beginning and end of the protocol. At $t=0$, a projective energy measurement is performed, then the system evolves under the time-dependent Hamiltonian, and at $t=\tau$ a second measurement is made. In this framework, work $W$ is a stochastic variable. For a single realization, it is given by $W[|m\rangle;|n\rangle]=E_{n}(\lambda(\tau))-E_{m}(\lambda(0))$, where $|m\rangle$ and $|n\rangle$ are the initial and final eigenstates with energies $E_{m}(\lambda(0))$ and $E_{n}(\lambda(\tau))$, respectively. The probability distribution of work is obtained by averaging over an ensemble of realizations esposito2009nonequilibrium
$\mathcal{P}(W)=\sum_{W=E_{n}-E_{m}}\delta\left(W-W[|m\rangle;|n\rangle]\right)\,p(|m\rangle\rightarrow|n\rangle),$ (55)
where $p(|m\rangle\rightarrow|n\rangle)$ denotes the probability of observing the transition $|m\rangle\rightarrow|n\rangle$. If the system is initially in a thermal Gibbs state $\rho_{eq}$, Eq. (55) reduces, as shown by Scandi et al. scandi2020quantum , to the standard TPM definition of work:
$\begin{split}&\mathcal{P}(W)\\ =&\sum_{W=E_{n}-E_{m}}\langle m(\lambda(0))|\rho_{eq}|m(\lambda(0))\rangle\,\big|\langle m(\lambda(0))|n(\lambda(\tau))\rangle\big|^{2}.\end{split}$ (56)
This expression represents the product of the probability of the system initially being in state $|m\rangle$ with the transition probability $|m\rangle\rightarrow|n\rangle$. However, since each projective energy measurement perturbs the system — and may even destroy its state — the full $\mathcal{P}(W)$ cannot be directly accessed in practice through this scheme scandi2020quantum ; deffner2019quantumthermo .
However, from the statistical theory, we know that any probability distribution can be reconstructed from its cumulant generating function (CGF) strasberg2022quantum , defined as
[…]
where $\lambda=\lambda(\tau)$, denotes the final value attained by the protocol. The CGF is additive under independent random measurements, and allows one to recover the probability distribution $\mathcal{P}(W)$ through an appropriate inverse Fourier transform. Moreover, all the statistical information of $\mathcal{P}(W)$, such as cumulants, can be computed by differentiating the CGF kardar2007book :
$\kappa_{W}^{(n)}=(-\beta)^{-n}\frac{d^{n}}{d\lambda^{n}}K(\lambda)\bigg|_{\lambda=0}.$ (58)
Since $W=\Delta F+W_{\text{diss}}$, the CGF naturally separates into two parts: a deterministic term depending only on the free-energy difference, and dissipative contribution due to the irreversible process, as shown by Scandi et al. scandi2020quantum
$K(\lambda)=-\beta\lambda\Delta F+K^{\text{diss}}(\lambda),$ (59)
[…]
Guarnieri et. al. guarnieri2024generalized , building on the elegant method of Shitara and Ueda shitara2016determining , established a landmark bridge between LRT and QFTs as
$K^{\text{diss}}(\lambda)=-\int_{0}^{\tau}dt\int_{0}^{\tau}dt^{\prime}\dot{\lambda}(t)\dot{\lambda}(t^{\prime})[g_{\lambda}*\Psi_{0}](t-t^{\prime}),$ (60)
where a function of $\lambda$, $g(\lambda)$, has to be convoluted with our old friend $\Psi_{0}$ to get the CGF of the dissipated work. The symmetry of $g(\lambda)=g(1-\lambda)$ ensures $K^{\text{diss}}(\lambda)=K^{\text{diss}}(1-\lambda)$ guarnieri2024generalized . Operating an inverse Laplace transform on it, gives our pre-acknowledged companion
$\frac{\mathcal{P}(W_{\text{diss}})}{\mathcal{P}(-W_{\text{diss}})}=e^{\beta W_{\text{diss}}},$ (61)
a special form of ESFT guarnieri2024generalized , now established in the quantum regime. The essential finding of Eq. (60) is that any phenomenological relaxation function can fully determine quantum work distribution in the linear regime. Moreover, this framework predicts all higher-order cumulants are positive, i.e., $\kappa_{W}^{(n)}\geq 0$ guarnieri2024generalized , leading to non-Gaussian tail (such as skewness and excess kurtosis), driven by quantum fluctuations scandi2020quantum . To quantify these non-classical effects, Guarnieri et al. guarnieri2024generalized derived a bound stronger than standard thermodynamic uncertainty relation (TUR) agarwalla2018assessing ; seifert2025stochastic . In LRT, the the conventional TUR reads (in terms of the Fano factor) guarnieri2024generalized :
$F_{W}=\frac{\sigma^{2}_{W_{diss}}}{\langle W_{diss}\rangle}\geq 2k_{B}T.$ (62)
While, Guarnieri et al. guarnieri2024generalized have rigorously proved that by defining a normalised probability distribution function $\widetilde{P}(\omega)$, associated with the system relaxation function $\widetilde{\Psi}_{0}(\omega)$, over the pseudomodes frequencies $\omega\in[0,\infty)$, the Fano factor can be expressed as:
$\begin{split}F_{W}&=~\langle\hbar\omega\coth(\beta\hbar\omega/2)\rangle_{\widetilde{P}}=2\langle\mathcal{E}_{\omega}\rangle_{\widetilde{P}}\\ &\geq~\hbar\langle\omega\rangle_{\widetilde{P}}\coth\!\left(\beta\hbar\langle\omega\rangle_{\widetilde{P}}/2\right)\end{split}$ (63)
Equation (63) provides a stronger bound than Eq. (62). Although one expects $F_{W}$ to saturate to classical behaviour in the long-time limit, $\mathcal{E}_{\omega}$ — the average energy of a quantum harmonic oscillator — includes the contribution of its average zero-point energy, which persists even in the vacuum state and prevents the system from becoming fully classical. At short times, however, vacuum fluctuations give rise to pronounced quantum signatures in the work distribution for various phenomenological Brownian-motion models.
V The way from here
“The first quantum revolution gave us new rules that govern physical reality. The second quantum revolution will take these rules and use them to develop new technologies.”
— Jonathan P. Dowling dowling2003quantum
Throughout this article, we have outlined a trajectory from 19th-century thermodynamics to modern formalisms, passing through the emergence of stochastic thermodynamics and culminating in the rise of quantum thermodynamics halpern2022quantum , formulated here within the language of LRT near equilibrium. In this framework, the QFT for TPM-based dissipated work reaches a height without relying on slow-driving or weak coupling aproximations guarnieri2024generalized . This perspective may shed light on non-Markovian effects and work fluctuations across phase transitions campbell2025roadmap , while remaining experimentally relevant since the Fano factor is directly measurable guarnieri2024generalized .
While the TPM definition is useful, quantum work is not an observable talkner2007fluctuation , and the approach suffers from serious drawbacks deffner2019quantumthermo . In particular, it neglects measurement back-action, and projective energy measurements are neither feasible nor practical deffner2019quantumthermo . These limitations motivate describing work fluctuations in terms of variations of physical quantities (VPQs) silva2025fundamental . A promising alternative is the two-time observable protocol, which is experimentally accessible in platforms such as trapped ions, plasma drops, superconducting qubits, and NMR setups silva2025fundamental . In this context, linear response theory may offer a natural framework for modeling these systems.
As our knowledge advances, we are steadily moving toward the ability to design highly efficient, optimised devices—such as quantum computers, sensors, and metrological tools — that harness quantum phenomena as fundamental resources. This heralds the onset of the second quantum revolution dowling2003quantum , in which quantum effects extend beyond atoms and molecules to profoundly shape technology, society, and civilisation.
VI Acknowledgements
We acknowledge IIT Kanpur for providing partial financial support.